A bond investor is considering two 10 year maturity bonds both rated A: the municipal bond is yielding 2.40% and the corporate bond is yielding 3.25%. At what marginal tax rate would the bond investor be indifferent between the two bonds? Enter your answer rounded off to two decimal points.
Answer #1

Municipal bond incomes are tax free. However the incomes on corporate bonds are taxable.

If T is the tax rate, the post tax yield on Corporate bond = 3.25% x (1 - T)

For indifference:

3.25% x (1 - T) = 2.40%

Hence, Marginal tax rate, T = 1 - 2.40% / 3.25% = 26.15%